This is a statistics problem, Can someone help? I am so confused.
An instructor gives a 100 point exam in which the grades are normally distributed. The mean is 60 and the standard deviation is 10. If there are 5% A's and 5% F's, 15% B's and 15% D's and 60% C, find the scores that divide the distribution into those categories

Answer by stanbon(75887)

An instructor gives a 100 point exam in which the grades are normally distributed. The mean is 60 and the standard deviation is 10. If there are 5% A's and 5% F's, 15% B's and 15% D's and 60% C, find the scores that divide the distribution into those categories.
-------------
The z-score below which 5% of the scores lie is -1.64485
The z-score above which 5% of the scores lie is +1.64485
-------------
Convert these z-scores to raw scores as follows:
x-60=-1.64485*10
x=60-16.4485
x=43.55 (this is the grade score below which 5% of the grades lie)
-------
x-60=.64485*10
x=60+16.4485
x=76.4485 (this is the grade score above which 5% of the grades lie)
--------
Follow the same procedure for the other percentages.
Let me know if you have trouble with this.
cheers,
Stan H.